# Stochastic Control Barrier Functions under State Estimation: From Euclidean Space to Lie Groups

### v1

@@ I-A Background and Motivation
Beyond uncertainty, an additional challenge stems from the geometry of the state manifold. In practice, the states of dynamical systems evolve on non-Euclidean spaces whenever rotational degrees of freedom are present, which has motivated the development of geometric control techniques for global property assessment, structural preservation, or vibrational stabilization (see, e.g., [37, 4, 7, 38]). From an application perspective, when the size of a robot is comparable to the scale of its environment, its shape becomes critical, particularly in narrow environments (see, e.g., [21]). For example, consider an unmanned aerial vehicle (UAV) maneuvering at high speed through a cluttered environment. In such settings, considering only the position of the UAV or approximating the UAV as a point mass is insufficient to design a safety filter, since collision avoidance depends on the UAV’s pose, which evolves on a Lie group. Furthermore, for stochastic systems, the importance of the geometry of the state manifold becomes more pronounced because it governs how uncertainty influences the state evolution (see, e.g., [5, 39, 28]), as illustrated in Fig. 1. This is different from the scenario where the state of a dynamical system is in Euclidean space and the noise is typically additive.
@@ I-B1 Safety for Stochastic Systems
In contrast, [11] proposes an online optimization-based control synthesis framework with finite-time safety guarantees for stochastic systems. However, it requires the CBF to be upper-bounded in order to construct probabilistic safety guarantees. Subsequently, [10] relaxes this requirement by assuming stronger structural information on the system (e.g., a bounded difference between the true and predictable CBF values), yielding tighter theoretical bounds. However, the control synthesis framework in [11, 10] does not account for the second-order moment information of CBFs so the resulting controllers do not adapt to the level of uncertainty. This may result in overly aggressive behavior when the level of uncertainty is relatively high, which will be demonstrated in Section IV-B. In addition, [29] proposed a probabilistic CBF (PCBF) framework, where the discrete-time CBF-based constraint is plugged into a chance constraint for stochastic systems, which is then converted into a computable deterministic surrogate constraint via probabilistic inequalities (e.g., Cantelli’s inequality) for control synthesis.
Another line of work is on safety verification. [32] proposes a barrier-certificate framework for stochastic systems with provable bounds on the probability of reaching unsafe sets. However, the barrier functions, the safe sets, and the system dynamics are assumed to be polynomial for tractable computation. Recently, [24, 25] investigates finite-time safety verification of stochastic systems by reducing the problem to deterministic verification on an eroded safe set constructed via Minkowski difference. However, even when an analytic function (e.g., a CBF) defining the original safe set is given, constructing such eroded sets may be challenging in complex environments or high-dimensional state spaces, similar to the difficulty encountered in free configuration-space construction in motion planning [21].
@@ II-B Lie Group
Let $G\in\mathcal{G}$ be a random variable on Lie group with its probability density function $f_{G}:\mathcal{G}\to\mathbb{R}_{\geq 0}$, then the mean $\mu=\mathbb{E}\!\left[\,G\,\right]\in\mathcal{G}$ is such that
@@ III Stochastic CBF under State Estimation Uncertainty
which are assumed to be sub-Gaussian, conditionally on $\mathcal{F}_{k}$, with variance proxies $\bar{\sigma}_{k+1}^{2}$ and $\bar{\tau}_{k}^{2}$, respectively. Based on Lemma II.2, (6) becomes
@@
$=\,$ $\widetilde{Y}_{k+1}-\mathbb{E}\!\left[\,Y_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right],$
and we have
$\mathbb{E}\!\left[\,\Delta_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right]=0.$
@@
Since $M_{k}$ and $\mathbb{E}\big[\,\widetilde{Y}_{k+1}\,\big|\,\mathcal{F}_{k}\,\big]$ are $\mathcal{F}_{k}$-measurable, then we have $\mathbb{E}\!\left[\,M_{k}\,\middle|\,\mathcal{F}_{k}\,\right]=M_{k}$, and
$=0.$
Hence, we have $\mathbb{E}\!\left[\,M_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right]=M_{k}$ a.s., $\forall k\in\mathbb{N}$, i.e., $(M_{k})_{k\geq 0}$ is a martingale. (Note that $\mathbb{E}\!\left[\,S_{1}\,\middle|\,\mathcal{F}_{0}\,\right]=S_{0}=0$, whereas $\mathbb{E}\!\left[\,M_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right]\neq 0$, $\forall k\geq 1$.)
@@
Since $\sum_{i=0}^{k-1}\alpha^{-(i+1)}\beta_{k}\sigma_{h_{k+1}}^{-}\geq 0$, we have
@@
Theorem III.1 provides an analytical, offline-computable bound of finite-time stochastic safety incorporating state estimation. Once the proxies $\bar{\sigma}_{k+1}^{2}$ and $\bar{\tau}_{k}^{2}$ are specified, the right-hand side of (17) can be computed without access to actual trajectories (i.e., without running experiments).
@@ IV-A Linear System with Affine CBF
Now that we have the closed-form expressions for $\bar{\sigma}_{k+1}^{2}$ and $\bar{\tau}_{k}^{2}$ which can be calculated offline via Riccati recursion as seen in (32) and (33), as discussed at the end of Section III we demonstrate the behavior of the bound in Theorem III.1 by applying the SEA-SCBF-QP (31) to a linear system with affine barrier function. Specifically, we consider (27) with $p_{k}\in\mathbb{R}^{2}$, $A=I$, $B=I\Delta t$, $H=I$, and $\epsilon_{k}\sim\mathcal{N}(0,0.03^{2}I)$, and (28) with $c=[0,1]^{\top}$ and b=0.5b=-0.5. We let $\alpha=1$, and $\beta_{k}=\beta\exp\!\left(-7\widetilde{Y}_{k}\right)$. To clearly illustrate safety violations in the Monte Carlo (MC) experiments, we set the level of some noise to be relatively higher than the system scale, and adopt an open-loop controller as the nominal control input that drives the state along the positive $x$-axis. Thus we let $\varepsilon_{k}\sim\mathcal{N}(0,\operatorname{diag}(0.03^{2},\sigma_{y}^{2}))$ to show the influence of the noise level in $y$-axis on the behavior of $P_{\text{\rm exit}}(T,x_{0})$. In addition, $\eta$ is obtained through (25).
@@ IV-B SCBF-based Motion Planning
The PCBF-based constraint is given by [29]
@@
A notable fact is that even though the SEA-PCBF-QP (40) typically results in more conservative actions, its safety rate is lower than the SEA-SCBF-QP (31). This is because when the estimated barrier function value is small, i.e., the estimated state is closer to an obstacle facet, the influence of the level of state estimation uncertainty has a greater impact in (31) than in (40). Especially when $\alpha$ is large and the level of the process noise is relatively higher than the one of the measurement noise, the effect of the state estimation uncertainty is suppressed in (39), resulting in relatively aggressive actions from (40) even when the state estimation uncertainty is high.
@@ V SEA-SCBF on Lie Groups
in which $g_{k},U_{k}\in\mathrm{SE}(3)$, $H:\mathrm{SE}(3)\times\mathbb{R}^{l}\to\mathcal{Z}$, $\varepsilon_{k}\sim\mathcal{N}(0,\Sigma_{\varepsilon_{k}})$, $\epsilon_{k}\sim\mathcal{N}(0,\Sigma_{\epsilon_{k}})$, $\Sigma_{\varepsilon_{k}}\in\mathbb{S}^{6}_{+}$, and $\Sigma_{\epsilon_{k}}\in\mathbb{S}^{l}_{+}$. The measurement manifold $\mathcal{Z}$ may be $\mathrm{SE}(3)$ itself or simply $\mathbb{R}^{3}$.
Then, we review state estimation with respect to (41) as follows (see, e.g., [3, 16, 8], for more detailed discussions on information fusion and state estimation on Lie groups). By using Lemma II.4, the prediction and update steps can be obtained as follows.

### v2

@@ I-A Background and Motivation
Beyond uncertainty, an additional challenge stems from the geometry of the state manifold. In practice, the states of dynamical systems evolve on non-Euclidean spaces whenever rotational degrees of freedom are present, which has motivated the development of geometric control techniques for global property assessment, structural preservation, or vibrational stabilization (see, e.g., [37, 4, 7, 38]). From an application perspective, when the size of a robot is comparable to the scale of its environment, its shape becomes critical, particularly in narrow environments (see, e.g., [21]). For example, consider an unmanned aerial vehicle (UAV) maneuvering at high speed through a cluttered environment. In such settings, considering only the position of the UAV or approximating the UAV as a point mass is insufficient to design a safety filter, since collision avoidance depends on the UAV’s pose, which evolves on a Lie group. Furthermore, for stochastic systems, the importance of the geometry of the state manifold becomes more pronounced because it governs how uncertainty influences the state evolution (see, e.g., [5, 39, 26, 28]), as illustrated in Fig. 1. This is different from the scenario where the state of a dynamical system is in Euclidean space and the noise is typically additive.
@@ I-B1 Safety for Stochastic Systems
In contrast, [11] proposes an online optimization-based control synthesis framework with finite-time safety guarantees for stochastic systems. However, it requires the CBF to be upper-bounded in order to construct probabilistic safety guarantees. Subsequently, [10] relaxes this requirement by assuming stronger structural information on the system (e.g., a bounded difference between the true and predictable CBF values), yielding tighter theoretical bounds. However, the control synthesis framework in [11, 10] does not account for the second-order moment information of CBFs so the resulting controllers do not adapt to the level of uncertainty. This may result in overly aggressive behavior when the level of uncertainty is relatively high, which will be demonstrated in Section IV-B. In addition, [29] proposes a probabilistic CBF framework, where the CBF-based constraint is reformulated as a chance constraint for stochastic systems, which is then converted into a computable deterministic surrogate constraint via probabilistic inequalities (e.g., Cantelli’s inequality) for control synthesis.
Another line of work is on safety verification. [32] proposes a barrier-certificate framework for stochastic systems with provable bounds on the probability of reaching unsafe sets. However, the barrier functions, the safe sets, and the system dynamics are assumed to be polynomial for tractable computation. Recently, [24, 25] investigates finite-time safety verification of stochastic systems by reducing the problem to deterministic verification on an eroded safe set constructed via Minkowski difference. However, even when an analytic function (e.g., a CBF) defining the original safe set is given, constructing such eroded sets may be challenging in complex environments or high-dimensional state spaces, similar to the difficulty encountered in analytic free configuration-space construction in motion planning [21].
@@ II-B Lie Group
Let $G\in\mathcal{G}$ be a random variable on a Lie group with its probability density function $f_{G}:\mathcal{G}\to\mathbb{R}_{\geq 0}$, then the mean $\mu=\mathbb{E}\!\left[\,G\,\right]\in\mathcal{G}$ is such that
@@ III Stochastic CBF under State Estimation Uncertainty
which are assumed to be sub-Gaussian, conditionally on $\mathcal{F}_{k}$, with variance proxies $\bar{\sigma}_{k+1}^{2}$ and $\bar{\tau}_{k}^{2}$, respectively. Based on Lemma II.2, we have
$=0.$
and (6) becomes
@@
$=\,$ $\widetilde{Y}_{k+1}-\mathbb{E}\!\left[\,Y_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right].$
@@
Since $M_{k}$ and $\mathbb{E}\big[\,\widetilde{Y}_{k+1}\,\big|\,\mathcal{F}_{k}\,\big]$ are $\mathcal{F}_{k}$-measurable, and $\mathbb{E}\!\left[\,\Delta_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right]=0$, then we have $\mathbb{E}\!\left[\,M_{k}\,\middle|\,\mathcal{F}_{k}\,\right]=M_{k}$, and $\mathbb{E}\!\left[\,M_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right]=M_{k}$ a.s., $\forall k\in\mathbb{N}$. Hence, $(M_{k})_{k\geq 0}$ is a martingale. (Note that $\mathbb{E}\!\left[\,M_{1}\,\middle|\,\mathcal{F}_{0}\,\right]=M_{0}=0$, whereas $\mathbb{E}\!\left[\,M_{k+1}\,\middle|\,\mathcal{F}_{k}\,\right]\neq 0$, $\forall k\geq 1$.)
@@
Since $\sum_{i=0}^{k-1}\alpha^{-(i+1)}\beta_{k}\sigma_{h_{k+1}}^{-}\geq 0$, then
@@
Theorem III.1 provides an offline-computable bound of finite-time stochastic safety incorporating state estimation uncertainty. Once the proxies $\bar{\sigma}_{k+1}^{2}$ and $\bar{\tau}_{k}^{2}$ are specified, the right-hand side of (17) can be computed without access to actual trajectories (i.e., without running experiments).
@@ IV-A Linear System with Affine CBF
Now that we have the closed-form expressions for $\bar{\sigma}_{k+1}^{2}$ and $\bar{\tau}_{k}^{2}$ which can be calculated offline via Riccati recursion as seen in (32) and (33), as discussed at the end of Section III we demonstrate the behavior of the bound in Theorem III.1 by applying the SEA-SCBF-QP (31) to a linear system with affine barrier function. Specifically, we consider (27) with $p_{k}\in\mathbb{R}^{2}$, $A=I$, $B=I\Delta t$, $H=I$, $\epsilon_{k}\sim\mathcal{N}(0,0.03^{2}I)$, $c=[0,1]^{\top}$, b=0.5b=-0.5, $\alpha=1$, and $\beta_{k}=\beta\exp\!\left(-7\widetilde{Y}_{k}\right)$. To clearly illustrate safety violations in the Monte Carlo (MC) experiments, we set the level of some noise to be relatively higher than the system scale, and adopt an open-loop controller as the nominal control input that drives the state along the positive $x$-axis. Thus we let $\varepsilon_{k}\sim\mathcal{N}(0,\operatorname{diag}(0.03^{2},\sigma_{y}^{2}))$ to show the influence of the noise level in $y$-axis on the behavior of $P_{\text{\rm exit}}(T,x_{0})$. In addition, $\eta$ is obtained through (25).
@@ IV-B SCBF-based Motion Planning
The PCBF-based chance constraint is given by [29]
@@
A notable fact is that even though the SEA-PCBF-QP (40) typically results in more conservative actions, its safety rate is lower than the SEA-SCBF-QP (31). This is because when the estimated barrier function value is small, i.e., the estimated state is closer to an obstacle facet, the influence of the level of state estimation uncertainty has a greater impact in (31) than in (40). Especially when the level of the measurement noise is relatively higher than that of the process noise, the effect of the state estimation uncertainty is suppressed in (39), resulting in relatively aggressive actions from (40) even when the state estimation uncertainty is high.
@@ V SEA-SCBF on Lie Groups
in which $g_{k},U_{k}\in\mathrm{SE}(3)$, $H:\mathrm{SE}(3)\times\mathbb{R}^{l}\to\mathcal{Z}$, $\varepsilon_{k}\sim\mathcal{N}(0,\Sigma_{\varepsilon_{k}})$, $\epsilon_{k}\sim\mathcal{N}(0,\Sigma_{\epsilon_{k}})$, $\Sigma_{\varepsilon_{k}}\in\mathbb{S}^{6}_{+}$, and $\Sigma_{\epsilon_{k}}\in\mathbb{S}^{l}_{+}$. The measurement manifold $\mathcal{Z}$ can be $\mathrm{SE}(3)$ itself or simply $\mathbb{R}^{3}$.
Then, we review state estimation with respect to (41) as follows (see, e.g., [8, 41, 16, 3], for more detailed discussions on information fusion and state estimation on Lie groups). By using Lemma II.4, the prediction and update steps can be obtained as follows.